OpenStudy (anonymous):

Solve the equation logbase27 x = 1 – logbase27 (x – 0.4).

OpenStudy (mertsj):

\[\log_{27}x=1-\log_{27}(x-.4) \]

OpenStudy (mertsj):

\[\log_{27}x+\log_{27} (x-.4)=1 \]

OpenStudy (mertsj):

\[\log_{27}x(x-.4) =1\]

OpenStudy (anonymous):

x^2-0.4x=27?

OpenStudy (mertsj):

\[27^{1}=x(x-.4)\]

OpenStudy (mertsj):

Solve that quadratic equation.

OpenStudy (anonymous):

Funny thing is I did all that xD

OpenStudy (anonymous):

I got stuck at the quadratic

OpenStudy (mertsj):

x^2-.4x=27 10x^2-4x-270=0

OpenStudy (mertsj):

\[2(x-5)(5x-27)=0\]

OpenStudy (mertsj):

And, of course, if you had trouble factoring, you could always use the quadratic formula.

OpenStudy (anonymous):

Nahh, I gots it from there thanks (=

OpenStudy (mertsj):

yw